# Fully Distributed Fog Load Balancing with Multi-Agent Reinforcement Learning

### v1

@@ III The Fog System Model
This section defines the system model of the Fog network and the formulation of the load balancing problem. Table II shows the definitions of the mathematical symbols used throughout this paper. We reproduce the realistic Fog environments used in our previous works [17, 18] to compare the proposed fully distributed multi-agent solution with the centralized single-agent solution in [17].
@@
$s_{t}$ The observed state at step $t$, from a set of possible states $S$
$a_{t}$ The agent’s action at step $t$, from a set of possible actions $A$
@@
The Fog environment has a mesh-like topology, with a central cloud, interconnected Fog nodes, and APs that work as gateways for their associated IoT devices. Load-balancing agents are deployed in these APs, where each agent makes its own decisions to distribute the load of its associated IoT devices between the Fog nodes in its collaboration region, i.e., its candidate Fog nodes. Agents in each collaboration region share a set of candidate Fog nodes, requiring the agents to optimize the resources of all candidate Fog nodes collaboratively. To be fully distributed, the agents must collaborate blindly, i.e., without sharing their experience or decision-making policies.
@@
The Fog network is defined by the set of $N=n_{1},n_{2},\cdots,n_{z}$ nodes described by their compute $IPT_{n}$ and memory $RAM_{n}$ resources [17, 18]. The Nodes, including $N^{AP}$ AP nodes, $N^{F}$ Fog nodes, and $N^{C}$ Cloud nodes, are connected through $L=l_{1},l_{2},\cdots,l_{z}$ bidirectional links. Each link is characterized by the pair ($n_{i}$, $n_{j}$) that it connects, its bandwidth $BW_{l}$, and its propagation delay $PR_{l}$. The network has a set of distributed applications $DA=da_{1},da_{2},\cdots,da_{z}$ that simultaneously run in the system; each application has a set of modules $M_{da}$ and dependency messages $D_{da}$ between these modules, i.e., workloads or requests.
@@
Nodes and network links have heterogeneous resources and must serve three different workload categories $|W|$ each with different compute requirements, i.e., light, medium, and heavy workloads. Workloads are generated as a Poisson Point Process using exponential distributions with three different scale parameters $\beta$ to simulate high, medium, and low workload generation rates, allowing for dynamic representation of job arrivals in our Fog environment. Nodes and links are modeled with M/M/1 queuing models to handle the generated workloads, giving the ability to provide insights into performance metrics such as queue lengths and waiting times. The agents must learn to minimize the number of queued jobs $\mathbb{Q}_{n}$ in the queue of each Fog node $n$ to minimize the workload waiting delay, reducing the end-to-end execution delay over the system [17].
To optimize this objective, each agent learns a policy to select the best Fog node for each generated workload. To do this, each agent in a collaboration region observes the number of jobs queued in every candidate Fog node in that region. These observations represent the state of the environment at each step, and it is the same for all agents in a collaboration region. This state is used to select an action $a$ from the set of possible actions $A$, corresponding to the set of candidate Fog nodes $N^{F}$ in that region. Each agent learns to select the action that maximizes its performance measure, i.e., the reward, which is evaluated at the end of every decision step. For each agent, the reward is the number of jobs queued in the Fog node selected in its previous action. To reduce the average system waiting delay, this number is negated to be maximized by the RL algorithm. This reduces the accumulation of workloads in Fog nodes, lowering overload probability and overall execution delay.
@@ IV Distributed RL Load Balancing
A fully independent Double Deep Q-Learning (DDQL) agent is deployed in each of the $N^{AP}$ APs of the Fog network. The state, action, and reward function of every agent at every decision step $t$ are defined as follows:
@@
  • State: $s_{t}=(w_{t},\mathbb{Q}_{n,t}:\forall n\in N^{CF})$, where $w$ represents the category of the workload to be assigned and $\mathbb{Q}_{n}$ represents the number of queued jobs in candidate Fog node $n$.
@@
  • Action: $a_{t}\in A$ is the decision to assign the current workload to the best candidate Fog node from the set of candidates considered for that agent.
@@
  • Reward: $r_{t}=-\mathbb{Q}_{a_{t-1},t}$ represents the number of queued jobs in the Fog node previously selected by that agent, i.e, the reward at decision step $t$ is the performance measure of the action taken at decision step $t-1$ (it is part of the observed state, and thus, does not need to be transmitted separately to each agent). The negation is used to represent the reward as a cost to be minimized instead of a utility to be maximized. The accumulation of these rewards is the expected discounted return $G=\sum_{t=0}^{\infty}\gamma^{t}r_{t}$, where $\gamma$ defines the importance of future rewards compared to recent ones.
In centralized single-agent solutions [2, 17], the state space consists of the number of queued jobs from all Fog nodes at each decision step, i.e., $\mathbb{Q}_{n,t}:\forall n\in N^{F}$. The centralized agent also requires the identity of the AP connected to the source IoT device, i.e., workload from $n^{AP}_{t}$ to be assigned at decision step $t$. When shifting to a distributed multi-agent solution in large-scale networks, considering all Fog nodes for every agent can unnecessarily increase the state and action spaces of each agent. However, it is sufficient to design the agents with feasible subsets of Fog nodes within a certain proximity. This leads to creating multiple groups of agents, each observing the queues of their own set of candidate Fog nodes $\mathbb{Q}_{n,t}:\forall n\in N^{CF}$.
In addition to increasing the state and action spaces, a centralized agent causes a significant delay for each workload assignment decision due to the three-way communication between the centralized agent and the workload AP (see Fig. 2LABEL:sub@fig:Topology_WF_C). For every generated workload in centralized load distribution solution, an assignment request is sent to the centralized agent along with the workload category and the identifier of its AP. After receiving this request, the centralized agent predicts the best Fog node for that workload using the provided information and the observed number of queued jobs in each Fog node in the system. The agent then responds to the requesting AP with the identifier of the selected Fog node. Finally, the AP assigns and sends the workload to the selected Fog node and waits for the next generated workload.
@@
With a distributed load balancing solution, however, each agent has this prediction capability locally, and hence, each agent makes its own assignment decisions (see Fig. 2LABEL:sub@fig:Topology_WF_D). As soon as the AP receives a workload, it predicts the best candidate Fog node for that workload by providing the workload category $w$ and the number of queued jobs in each candidate Fog node. The workload is then immediately assigned and transmitted to the selected Fog node, minimizing the decision delay compared to a centralized solution.
In addition to the decision-making delay induced by the agent’s location, i.e., centralized vs. distributed, observing the number of queued jobs in each Fog node causes a significant delay, often neglected in the literature. In a centralized solution, each Fog node sends its queue information to the centralized agent as in Fig. 3LABEL:sub@fig:Topology_info_C. In a distributed solution, however, each agent/AP needs to receive this information from all candidate Fog nodes (see Fig. 3LABEL:sub@fig:Topology_info_D). To do this in a real-world environment, Fog nodes need to multi-cast their queue information, introducing a significant delay compared to a centralized solution.
@@
Frequently multi-casting the observations, i.e., for every generated workload, in large-scale networks consumes a significant amount of network bandwidth. Hence, providing this information in real-time for every workload assignment decision becomes infeasible, especially with a large number of Fog nodes and IoT devices, and high workload generation rates. Therefore, addressing the frequency of transferring this information between the agents becomes essential. In addition, dividing large-scale networks into smaller subsets of collaboration regions limits the transfer of this information to fewer agents within a smaller geographical scope, thereby reducing the communication overhead.
A distributed load balancing solution allows the network to be clustered into smaller, possibly overlapping, regions based on geographical proximity between Fog nodes and APs. This significantly reduces the state and action spaces by only considering feasible Fog nodes for the agents in each collaboration region, which helps reduce the computational complexity of the solution. The agents in each region will then collaborate within their respective boundaries to optimize the resources of their feasible candidate Fog nodes. Unlike centralized load balancing solutions, where all Fog nodes must be included in the observation of the centralized agent, whether they are feasible options for all APs or not.
In this paper, we evaluate our fully distributed solution on a single collaboration region, where all agents collaborate to optimize the resources of the candidate Fog nodes in their region. To show the scalability of this solution to a larger network with multiple collaboration regions, these agents must be able to independently collaborate to minimize the number of queued jobs in each candidate Fog node. To learn independently, the agents must collaborate blindly, i.e., without knowing the policies and actions of the other agents in the system.
@@
Given two non-overlapping sets of candidate Fog nodes $\mathcal{F}_{1}\cap\mathcal{F}_{2}=\emptyset$ in two collaboration regions. Each region is managed by a group of agents that blindly collaborate to optimize the resources of their candidate Fog nodes without knowing the actions of other agents in the system (see Fig. 4). If the two regions share a subset of candidate Fog nodes $\mathcal{F}_{1}\cap\mathcal{F}_{2}=\mathcal{F}_{S}$, the agents of each region can still blindly collaborate to optimize the resources of their set of candidate Fog nodes and the shared set of candidate Fog nodes.
@@
Assuming the agents in an isolated region can efficiently manage the resources of their candidate Fog nodes by blindly collaborating without the need to observe the actions, nor the mechanism of making those actions, of other agents in their collaboration group. This assumption holds for any number of isolated regions each with their own set of candidate Fog nodes. Now, assume that a region $\mathcal{R}_{1}$ with $\mathcal{F}_{1}$ candidate Fog nodes share $\mathcal{F}_{S}$ candidate Fog nodes with region $\mathcal{R}_{2}$ (see Fig. 4). Suppose there exists an agent in $\mathcal{R}_{2}$ that does not collaborate with the agents in its group leading to inefficient management of the Fog resources in $\mathcal{F}_{2}$, including those in $\mathcal{F}_{S}$. This implies that there is a breakdown in the collaborative optimization process. However, this contradicts our initial assumption that within each region, the agents efficiently collaborate to optimize the resources of their candidate Fog nodes. Therefore, we conclude that the agents of both regions will collaborate to optimize the shared resources in $\mathcal{F}_{S}$ as well as those of their own. Therefore, the lemma holds. ∎
@@
In this work, we provide empirical evidence under a realistic setup supporting the ability of our independent distributed RL agents to optimally distribute the load within a collaboration region. Each agent acts according to the current environment state, and the learned policy $\pi$. In real-world environments, it is impractical to observe the state of every candidate Fog node for every workload assignment decision due to the network delay, which is intolerable for real-time applications. Observing the state of all Fog nodes in large-scale networks necessitates waiting for the state of the last node to transmit its load information to complete the environment observation. Hence, the effect of the observation delay is impacted by the scale of the Fog system, the number of IoT devices, and the rate of workload generation. In the literature, however, observations are always assumed to be immediately available for the agent at each decision step [19], rendering those solutions impractical for real-world implementations.
To provide a practical solution with real-world constraints, we use a realistic approach to observe the number of jobs queued in each Fog node. In this approach, the agents collect observations using interval-based multi-casting protocols, i.e., sending queue information on predetermined intervals. Using the same observation for multiple decision steps will lead to sub-optimal solutions, depending on the frequency of collecting the observations. Hence, the smallest practical interval for the multi-casting protocol must be considered to transfer the queue information to all agents. In this work, for instance, Fog nodes transfer their queue information every 3 seconds, which is a typical interval for Gossip [20] multi-casting protocols.
@@
After learning the optimal policy, we allow the agents to stop training, and use the learned models for inference only. This mitigates the need for unnecessary computations and leverages the simple arithmetic and array operations of lightweight inference models. The agents can then detect changes in the environment by measuring degradation in performance, which is used to trigger the fine-tuning process for their models to adapt to the new environment. Fine-tuning the learned models in a lifelong learning process can speed up convergence, improve data efficiency, and provide better generalization than learning from scratch [23]. In this work, for instance, the agents experience three significant surges in workload generation rates, causing three inference/fine-tuning epochs. Fine-tuning can also be used to initialize new agents by leveraging the expertise of existing trained agents in the network (see Fig. 5).
@@
Exploiting previously learned knowledge $\mathcal{I}_{\mathcal{S}}$ from source task $\mathcal{S}$ to improve the learning on a related target task $\mathcal{T}$ requires the tasks to have a certain level of similarity [24]. Hence, the learned policy on the target environment $E_{\mathcal{T}}$ depends on the information from both source and target tasks, i.e., $\pi_{\theta}=\phi(\mathcal{I}_{\mathcal{S}}{\sim}E_{\mathcal{S}},\mathcal{I}_{\mathcal{T}}{\sim}E_{\mathcal{T}}):S^{\mathcal{T}}\rightarrow A^{\mathcal{T}}$. This information includes the parameters of the Q network ( $\theta_{\mathcal{T}}\leftarrow\theta_{\mathcal{S}}$) and the $x$ recent interactions of the replay buffer ($\mathcal{B}_{\mathcal{T}}\leftarrow\text{RecentExperiences}(\mathcal{B}_{\mathcal{S}},x)$).

### v2

@@ III The Fog System Model
This section defines the system model of the Fog network and the formulation of the load balancing problem. Table II shows the definitions of the mathematical symbols used throughout this paper. We reproduce the realistic Fog environments used in our previous works [17, 18] to compare the proposed fully distributed (multi-agent) solution with the centralized (single agent) solution in [17].
@@
$s_{t}$ State at step $t$, from a set of possible states $S$
$a_{t}$ Action at step $t$, from a set of possible actions $A$
@@
The Fog environment has a mesh-like topology, with a central cloud, interconnected Fog nodes, and APs that work as gateways for their associated IoT devices. Load-balancing agents are deployed in these APs, where each agent makes its own decisions to distribute the load of its associated IoT devices between the Fog nodes in its collaboration region, i.e., its candidate Fog nodes. A group of agents in each collaboration region manages a set of candidate Fog nodes, requiring the agents to jointly optimize the resources of all candidate Fog nodes in that region. To be fully distributed, the agents must collaborate blindly without exchanging any kind of coordination information, like experiences, actions, or policies parameters.
@@
In our simulated environments, the Fog network is defined by the set of $N=n_{1},n_{2},\cdots,n_{z}$ nodes described by their compute $IPT_{n}$ and memory $RAM_{n}$ resources [17, 18]. The Nodes, including $N^{AP}$ AP nodes, $N^{F}$ Fog nodes, and $N^{C}$ Cloud nodes, are connected through $L=l_{1},l_{2},\cdots,l_{z}$ bidirectional links. Each link is characterized by the pair of nodes ($n_{i}$, $n_{j}$) that it connects, its bandwidth $BW_{l}$, and its propagation delay $PR_{l}$. A set of distributed applications $DA=da_{1},da_{2},\cdots,da_{z}$ simultaneously run in the system; each application has a set of modules $M_{da}$ and dependency messages $D_{da}$ between these modules, i.e., workloads or requests.
@@
Nodes and network links have heterogeneous resources and must serve different workload categories $|W|$ each with different compute requirements, i.e., light, medium, and heavy workloads. Workloads are generated as a Poisson Point Process using exponential distributions with three different scale parameters $\beta$ to simulate high, medium, and low workload generation rates. This allows for a dynamic representation of job arrivals in our Fog environment, where small $\beta$ values increase the probability of having smaller time intervals between subsequent workloads, leading to higher generation rates. Nodes and links are modeled with M/M/1 queuing models to handle the generated workloads, giving the ability to provide insights into performance metrics such as queue lengths and waiting times. To minimize the waiting time before being served by Fog nodes, the agents must learn to reduce the number of queued jobs $\mathbb{Q}_{n}$ in the queue of each Fog node $n$. Reducing workload accumulation in each Fog node reduces Fog overload probability and the end-to-end execution delay of IoT application loops [17].
To optimize this objective by independent agents, the agents need to learn independent policies that can balance the generated load between Fog nodes. To do that without exchanging coordination information, the agents observe the number of jobs queued in each Fog node in their region (as environment state $s$) to select the Fog node with the smallest load (as action $a$). Each agent takes its actions, from the set of possible actions $A$ corresponding to the set of candidate Fog nodes $N^{F}$ in its region, independently from all other agents. To reduce waiting time collaboratively by all agents, they learn to minimize the common objective of reducing workload accumulation in all Fog nodes. Therefore, the reward $r$ that evaluates their most recent action is defined by the load in the most recently assigned Fog node by that action, encouraging the agents to select the least loaded Fog nodes collaboratively. This rewarding mechanism allows the agents to jointly reduce workload accumulation in the Fog nodes of their region without sharing any collaboration information.
@@ IV Distributed RL Load Balancing
A fully independent Double Deep Q-Learning (DDQL) agent is deployed in each of the $N^{AP}$ APs of the Fog network. The agents in a given collaboration region are designed to manage the set of candidate Fog nodes $N^{F_{C}}$, i.e., a subset of Fog nodes from the larger network, in that region only. Hence, the size of state and action spaces can vary in different regions based on the number of candidate Fog nodes in each region, giving flexibility in forming those regions based on the needs of Fog service providers. This limits state and action spaces, and immediate rewards, to be collected from a smaller number of nearby Fog nodes, reducing the overhead of observing Fog queue information before every workload assignment decision. Hence, the state, action, and reward function of every agent in a given region, at every decision step $t$, are defined as follows:
@@
  • State: $s_{t}=(w_{t},\mathbb{Q}_{n,t}:\forall n\in N^{F_{C}})$, where $w$ is the number of compute instructions required by the workload to be assigned (representing its category) and $\mathbb{Q}_{n}$ is the number of queued jobs in each candidate Fog node $n$.
@@
  • Action: $a_{t}\in A$ is the decision to assign the current workload to a candidate Fog node from the set of candidate Fog nodes in the region of that agent.
@@
  • Reward: $r_{t}=-\mathbb{Q}_{a_{t-1},t}$ is the number of queued jobs in the Fog node previously selected by that agent, i.e, the reward at decision step $t$ is the performance measure of the action taken at decision step $t-1$. Because the value of the reward is part of the observed state, it does not need to be transmitted separately to each agent. This value is then negated to represent the reward as a cost to be minimized instead of a utility to be maximized. The accumulation of these rewards is the expected discounted return $G=\sum_{t=0}^{\infty}\gamma^{t}r_{t}$, where $\gamma$ defines the importance of future rewards compared to recent ones. Giving importance to future rewards encourages the agents to learn a load distribution strategy that guarantees minimizing load accumulation in all Fog nodes over the long term.
Our reward function enables the distribute agents to independently learn different load distribution strategies that keep the load in all Fog nodes minimal at all times. Each agent can discover its best strategy by considering its nearby Fog nodes before selecting remote Fog nodes in its region; this can be independently learned faster than learning a common policy for all agents. Hence, these independent agents can converge faster than single-agent solutions, as well as multi-agent solutions that lead to a common policy, i.e., explicitly and implicitly learn a common policy with CTDE and parameter/experience sharing, respectively. Also, learning independent policies allows us to reduce the state and action spaces by managing smaller subsets of Fog nodes, forming a collaboration region for each group of agents. Moreover, distributed solutions can provide near real-time decisions for every generated workload by avoiding the three-way communication overhead of centralized solutions (see Fig. 2LABEL:sub@fig:Topology_WF_C).
@@
In centralized solutions, a request for every generated workload is sent to a distanced agent with the category of the workload and the identifier of its AP. The agent notifies the AP of the suitable Fog node to serve that workload based on workload category, AP identifier, and the load in each Fog node in the network. Only then, the AP can send the workload to the selected Fog node. With distributed solutions, however, agents make their own assignment decisions locally (see Fig. 2LABEL:sub@fig:Topology_WF_D). The workload is then immediately sent to the selected Fog node, minimizing the decision delay compared to a centralized solution.
The time needed to observe Fog load information and the associated overhead of this process is often neglected in the literature [19]. With a single-agent solution, all Fog nodes send their load information to that agent for every generated workload, as in Fig. 3LABEL:sub@fig:Topology_info_C. Although our distributed agents are grouped to manage smaller subsets of Fog nodes in proximity, this overhead is still significant since every agent needs to collect this information from every candidate Fog node in their region (see Fig. 3LABEL:sub@fig:Topology_info_D). This makes MARL observation communication overhead, for every decision, exponentially proportional to the number of Fog nodes and agents in each region.
@@
Besides consuming bandwidth, observations can not be readily available before every decision due to communication delays, especially in large-scale networks with high generation rates. A more practical approach for deployment-ready solutions is to multi-cast such information at predefined intervals. This might lead to sub-optimal decisions due to using the same observation to make multiple assignment decisions, creating a non-stationary environment for the independent agents due to these outdated observations. This problem does not exist with centralized and coordination-based RL because the new environment state depends on a common policy behavior, not on the joint (unobserved) actions of independent agents. Being unrealistic to observe the real-time load on each Fog node before every decision, makes it more practical to consider a single observation for several actions within a given interval. Increasing observation frequency, using the smallest interval of existing multi-casting protocols (3 seconds for Gossip [20]), reduces the impact of non-stationarity.
In addition, dividing large-scale networks into smaller subsets of collaboration regions limits the transfer of Fog load information to fewer agents within a smaller geographical scope, thereby reducing the communication overhead. These regions can be designed based on the requirements of Fog service providers with the possibility of creating overlapped regions by sharing Fog resources between them. Managing fewer Fog nodes in proximity reduces state and action spaces for each agent and avoids the need to manage infeasible remote Fog nodes in large-scale networks. To ensure optimal load distribution in overlapped regions, the agents must be able to blindly learn independent load distribution strategies without knowing the policies and actions of the other agents. By independently learning to select the least loaded Fog node, our agents minimize the load on all Fog nodes in their regions, including shared Fog nodes.
@@
Given two non-overlapping sets of candidate Fog nodes $\mathcal{F}_{1}\cap\mathcal{F}_{2}=\emptyset$ in two collaboration regions (see Fig. 4). The resources in each region are optimized by a group of agents that blindly collaborate without knowing the actions of other agents in that region. If both regions share a subset of candidate Fog nodes $\mathcal{F}_{1}\cap\mathcal{F}_{2}=\mathcal{F}_{S}$, the agents of both regions can still blindly collaborate to optimize local and shared Fog resources.
@@
Assuming the agents in an isolated region can blindly optimize their isolated Fog resources without knowing the actions of the other agents in their region, nor the mechanism of making those actions. Now, assume that a region with $\mathcal{F}_{1}$ candidate Fog nodes share $\mathcal{F}_{S}$ candidate Fog nodes with a region having $\mathcal{F}_{2}$ candidate Fog nodes (see Fig. 4). Suppose an agent in one of the regions does not collaborate with the other agents, leading to inefficient utilization of Fog resources in its region, including $\mathcal{F}_{S}$ resources. This implies that there is a breakdown in the collaborative optimization process. However, this contradicts our initial assumption that within each region, the agents can blindly optimize the resources of their candidate Fog nodes. Therefore, we conclude that the agents of both regions will collaborate to optimize both local and shared resources. Thus, the lemma holds. ∎
@@
This paper provides empirical proof supporting the ability of our agents to blindly learn independent load distribution strategies that will jointly optimize the Fog resources in their isolated region without any coordination between these agents. Validating the efficiency of our fully distributed solution under a realistic setup of one isolated region proves the solution’s applicability in a global-scale multi-region network with isolated and overlapped collaboration regions. A fully distributed solution requires each agent to independently learn an efficient load distribution strategy, i.e., policy $\pi$, based on the category of the workload to be assigned and the current load in each Fog node in the region. Learning independent policies to optimize a common objective, i.e., selecting the least loaded Fog node, blindly encourages each agent to choose nearby Fog nodes before considering distanced ones.
@@
To mitigate the computational cost of continuous training in a practical deployment, each agent adopts our lifelong learning framework previously proposed in [2] (see Fig. 5). This framework can be used to initialize new agents in a simulation, or by leveraging the expertise of existing agents in the network, before fine-tuning them in their intended deployment environment, mitigating the impact of bad actions when training from scratch in real networks. With this framework, the agents can stop training when they converge to the optimal policy (performance saturation) and only retrain in case of significant performance degradation due to environmental changes. When convergence is detected, i.e., when the average reward does not improve further, the agents extract lightweight inference models to provide inference using only simple arithmetic and array operations.
@@
Changes in the environment can then be detected by observing a significant degradation in the average reward, i.e., an increase in the average load accumulation in the region, triggering the need for agents to perform a corrective measure to adapt to those changes. The agents are then fine-tuned using transfer learning to speed up convergence, improve data efficiency, and provide better generalization than learning from scratch [23]. Hence, policy network parameters are transferred from source to target task ($\theta_{\mathcal{T}}\leftarrow\theta_{\mathcal{S}}$) along with $x$ recent replay buffer interactions ($\mathcal{B}_{\mathcal{T}}\leftarrow\text{RecentExperiences}(\mathcal{B}_{\mathcal{S}},x)$), providing fast and efficient lifelong adaptation to changes [2]. Since effective transfer learning requires source and target tasks to share a certain level of similarity [24], we introduce changes only through significant surges in workload generation rates. Using transfer learning allows the agents to learn robust policies using information $\mathcal{I}$ from both source and target environments, i.e., $\pi_{\theta}=\phi(\mathcal{I}_{\mathcal{S}}{\sim}E_{\mathcal{S}},\mathcal{I}_{\mathcal{T}}{\sim}E_{\mathcal{T}}):S^{\mathcal{T}}\rightarrow A^{\mathcal{T}}$.